# Compton scattering from superstrings

### v1

@@ 1 Introduction
An important insight arising from studies of quantum scattering amplitudes is that gravitational amplitudes are simpler than one would expect [1]. The double copy prescription determines that the gravitational amplitudes are obtained as a product of two Yang-Mills amplitudes [2, 3, 4]. The original formulation by Kawai, Lewellen and Tye states that a closed string amplitude at the tree level is given by a sum over the products of two open string amplitudes at tree level [5]. The current trend is invaluable for the physics of gravitational waves [6, 7, 8, 9] due to the birth of gravitational wave astronomy, driven by landmark observations made by the LIGO and Virgo collaborations [10, 11, 12, 13, 14].
@@
In this context, we have witnessed a particularly intense flurry of activities about the gravitational interaction of spinning particles [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60]  [^4] For a much larger body of older and current research please see the references within [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60].. Indeed, gravitational dynamics of spinning objects has been a subject of vigorous study for several years [61, 62, 63, 64]. An outstanding problem emerging in this regard is the evaluation of the gravitational Compton scattering amplitude – the opposite-helicity version is a key ingredient in the evaluation of the conservative potential of compact binary systems with spin [41, 42, 43, 65, 66] and also waveform calculations with spin [60, 59, 67]. However, previous calculations in the literature displayed the presence of non-physical poles for massive particles with quantum spins greater than one, in the electromagnetic case, and greater than two, in the gravitational case [68, 69]  [^5] More precisely, the opposite-helicity Compton amplitude built from BCFW recursion relations have unphysical poles because this procedure is no longer predictive in the case of higher spins. This is due to the fact that the product of BCFW shifted three-point amplitudes does not approach zero fast enough as the shift parameter goes to infinity [25]. Nevertheless, we remark that if one starts with a local Lagrangian the amplitudes calculated from it using standard Feynman rules are expected to be manifestly local; see Ref. [41] for a calculation for the case of massive spin 3 electromagnetic Compton.. Important ongoing research is being conducted to address this problem [28, 25, 21, 22, 27, 29, 30, 31, 32, 33, 58].
This scenario opens up a great opportunity to investigate string amplitudes. It is known that one of the most notable features of string theory is the presence of an infinite tower of massive spinning modes in its spectrum required for consistency [70, 71]. However, some difficulties have been highlighted in the literature [72, 73, 42]; in particular, the classical limit of 3-point string amplitudes associated with leading Regge states of the open and closed superstring differ from the expected result for a Kerr black hole. This seems to indicate that using string theory to describe Kerr black holes is at the very least a formidable and challenging task.
@@
In this paper we wish to report significant progress. As an unexpected result we extract from a string-theory based calculation the correct tree-level Compton amplitude for a Kerr black hole for quantum spin two – the analogous result for $\sqrt{\textrm{Kerr}}$ is for quantum spin one, which we also recover here  [^6] Paolo di Vecchia and Francesco Alessio have also obtained this result through a slightly different calculation. We thank them for sharing their notes with us.. Having succeeded in recovering this important benchmark, we propose a candidate Compton amplitude which is free from any spurious poles and valid to any integer quantum spin. In order to obtain such amplitudes we impose three plausible physical constraints, to be discussed later on in this paper. With these simple rules we manage to obtain sensible field-theory Compton amplitudes which are physically sound. We are employing natural units, with $\hbar=c=1$.
@@ 2.1 Open strings
Let us calculate the Compton amplitude involving two photons and two massive spin-$S$ particles. For simplicity in this paper we will only study the NS sector of the string. We consider the Minkowski metric as $\eta_{\mu\nu}=\textrm{diag}(-,+,+,\cdots,+)$ and Mandelstam variables defined as $s_{ij}=-(k_{i}+k_{j})^{2}$, with $s_{12}=s$, $s_{13}=t$ and $s_{14}=u$. In addition, $k_{1}^{2}=k_{3}^{2}=-m^{2}$. Here $s+t+u=2m^{2}=2(S-1)/\alpha^{\prime}$, and the leading Regge state satisfies $m^{2}\alpha^{\prime}=n=S-1$. The vertex operators required here are the following [71]
@@
We employ a given representation of the massive polarization tensors $\epsilon_{(S)}$ (they are symmetric, traceless and transverse in all indices) of spin-S particles in which they are written as symmetric tensor products of massive spin-$1$ polarization vectors $\epsilon_{1}$, namely $\epsilon_{(S)}=\bigotimes_{S}\epsilon_{1}$. We should also consider $V_{0}^{(0)}$ – in order to cancel the background ghost charge on a genus $g$ Riemann surface, a $g$-loop superstring amplitude needs an overall superghost charge of $22g-$ 2. Finally, we will take $V_{0}^{(0)}$ to represent the photon operator in the zero picture. Our disk amplitude on the leading Regge trajectory reads
@@
where the factor $e^{-\lambda}$ comes from the Euler-number term in the action. For convenience we write the amplitude as
@@
where the dot products refer to the contraction of the higher-spin indices for legs 1 and 3. Henceforth we are suppressing the higher-spin indices for clarity. Moreover, we consider $z_{1}>z_{2}>z_{4}$. This means that we integrate $z_{3}$ over $[z_{4},z_{2}]$ to ensure cyclic ordering $(1,2,3,4)$ of the insertion points. Here we take $z_{1}\to\infty$, $z_{2}=1$ and $z_{4}=0$, which implies that $z_{3}\in[0,1]$.
@@ 2.2 Closed strings
$M_{4}({\bf 1}^{S},2^{h_{2}\tilde{h}_{2}},{\bf 3}^{S},4^{h_{4}\tilde{h}_{4}})$ $=$ $(-1)^{n}\left(\frac{\pi\alpha^{\prime}}{2}\right)^{-1}\sin\left(\frac{\pi\alpha^{\prime}s}{4}\right)$ (8)
@@
$M^{\alpha\gamma,\beta\delta}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $(-1)^{n}\left(\frac{\pi\alpha^{\prime}}{2}\right)^{-1}\sin\left(\frac{\pi\alpha^{\prime}s}{4}\right)\sin\left(\frac{\pi\alpha^{\prime}u}{4}\right)$ (9)
@@ 3 Electromagnetic Compton amplitudes for all spins
Our aim is to extract from the above string results terms potentially relevant for the classical description of spinning compact binary systems. Superstring amplitudes are famously defined in ten dimensions, so we need to perform a compactification down to four dimensions; we follow the prescription given by Ref. [72] and define suitable four-dimensional polarizations and momenta. Moreover, we also isolate the poles corresponding to the mass level $n$; these are connected to the s-channel and the u-channel. To extract the associated field-theory amplitudes for generic spins, we consider both momenta associated with the external photons as being soft, $q\to\hbar q$. Observe that keeping only the level $n$ poles corresponds to the leading 1/$1/\hbar$ term in the soft expansion. Then at leading order in a soft expansion we find
$A^{\alpha\beta}_{t}\Bigg{|}_{\textrm{field theory}}\equiv{\cal A}^{\alpha\beta}_{S}=\frac{1}{[\Gamma(n)]^{2}}\frac{{\cal A}_{0}^{\alpha\beta}}{(s\alpha^{\prime}-n)(u\alpha^{\prime}-n)}+{\cal O}(\alpha^{\prime}t).$ (10)
where the subscript $S$ in ${\cal A}^{\alpha\beta}_{S}$ denotes the spin of the massive particle. Before continuing the discussion for generic spins, let us study the lowest-spin massive boson. The polarization tensors are given by $\epsilon_{a_{1}a_{S}}({\bf 1})=\epsilon_{a_{1}}({\bf 1})\epsilon_{a_{S}}({\bf 1})$ and $\epsilon_{c_{1}c_{S}}({\bf 3})=\epsilon_{c_{1}}({\bf 3})\epsilon_{c_{S}}({\bf 3})$. By taking a close look at the corresponding expression, one quickly finds terms that describe the Compton amplitude for scattering of massive spin $1$ particles. To separate such terms from the rest we consider a suitable compactification of the polarization vectors $\epsilon_{a_{1}}({\bf 1})$ and $\epsilon_{c_{1}}({\bf 3})$ – we take these to be in the “fifth dimension” (such as $\epsilon^{d=10}({\bf 1})=(0,0,0,0,1,{\bf 0})$). After performing this compactification, contraction with the polarizations, and a few algebraic manipulations, we arrive at the following expression of the Compton amplitude describing the interaction between photons and massive spin-$1$ particles:
@@
where $h_{2},h_{4}$ are the helicities of the photons and $F_{i}^{\mu\nu}=2\epsilon^{[\mu}(i)q_{i}^{\nu]}$. To obtain this expression we wrote $\epsilon_{a_{S}}({\bf 1})\epsilon_{c_{S}}({\bf 3})={\cal J}_{a_{S}c_{S}}/2+\epsilon_{(a_{S}}({\bf 1})\epsilon_{c_{S})}({\bf 3})$, where we have defined ${\cal J}^{\alpha\gamma}\equiv-i\epsilon^{\mu}({\bf 1})(M^{\alpha\gamma})_{\mu\nu}\epsilon^{\nu}({\bf 3})$, $(M^{ab})_{c}^{\ d}=2i\delta^{[a}_{\ c}\eta^{b]d}$ being the Lorentz generator in the spin-$1$ representation. Furthermore:
@@
Up to a numerical overall factor, the result (11) perfectly matches the one previously derived by others [22] (except for possible sign differences due to different conventions on momenta). By rewriting this expression with the help of spinor-helicity methods, we also find that this amplitude agrees with other known expressions in the literature [68, 69].
Observe that $\alpha^{\prime}$ is effectively parameterizing the expansion in the soft photon momenta. Nevertheless, to see that the approach of taking directly $\alpha^{\prime}\to 0$ is not, in general, the proper way to obtain the field theory limit which is useful in the present context, let us quote the amplitude for the spin-2 leading Regge state. The calculation is rather involved but straightforward. After the dust settles, we find that
@@
where ${\cal J}^{ab}\equiv-\frac{i}{2}\epsilon^{c_{1}c_{2}}({\bf 1}){(M^{ab})_{c_{1}c_{2}}}^{d_{1}d_{2}}\epsilon_{d_{1}d_{2}}({\bf 3})$, ${(M^{ab})_{c(2)}}^{d(2)}=4i\delta_{(c_{1}}^{[a}\eta^{b](d_{1}}\delta_{c_{2})}^{d_{2})}$ being the Lorentz generator in the spin-$2$ representation. This amplitude in the classical limit will produce contributions up to the quadrupole term. However, for a spin-$S$ quantum particle, the classical counterpart should produce an expansion up to the $2S$-th power of the classical spin. Besides, strictly speaking, the $\alpha^{\prime}\to 0$ amounts to projecting out all massive spinning particles from the spectrum. These are clearly indications that the limit $\alpha^{\prime}\to 0$ is not adequate to our purposes.
Given the precedent successful computation, Eq. (11), let us proceed to the general case. Following previous prescriptions, we consider the compactification of two massive polarization vectors, namely $\epsilon_{a_{n}}({\bf 1})$ and $\epsilon_{c_{n}}({\bf 3})$. Furthermore, we consider only terms with $k=l$ in the double summations in the full expression for ${\cal A}_{0}^{\alpha\beta}$ (see appendix) as these are the ones that have explicit inverse propagators in the numerators. Our candidate electromagnetic Compton amplitude takes the form

### v2

@@ 1 Introduction
An important insight arising from contemporary studies of quantum scattering amplitudes is that gravitational amplitudes are simpler than one would expect from a brute-force Feynman diagram calculation [1]. The double copy prescription determines that the gravitational amplitudes are obtained as a product of two Yang-Mills amplitudes. The original formulation by Kawai, Lewellen and Tye states that a closed string amplitude at the tree level is given by a sum over the products of two open string amplitudes at tree level [2]. More recently, Bern, Carrasco and Johansson demonstrated that gravitational numerators are, in a sense, the square of the kinematic numerators of the Yang-Mills theory [3, 4, 5, 6], which has revealed another aspect of the double copy. This approach also connects classical solutions of Yang-Mills theory and gravity. In particular, point charges in Yang-Mills theory map to point sources in gravity [7, 8, 9, 10]. There are also indications that the double copy may encompass bound states and particles with spin [11, 12], and should also be relevant for higher-derivative gravity theories [13, 14, 15, 16, 17, 18, 19].
This current trend can be expected to be of extreme importance for the physics of gravitational waves. Indeed, the broader interest in this program is to scrutinize the application of modern scattering amplitude techniques to address the two-body gravity problem. This question arises naturally due to the birth of gravitational wave astronomy, driven by landmark observations made by the LIGO and Virgo collaborations [20, 21, 22, 23, 24]. Such techniques complement several methods presented in the literature and have been of utmost significance in fostering interdisciplinary collaborations, including the exploration of new research directions. Amplitudes have already been successfully applied to understand aspects of the general relativistic two-body problem [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 6, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78]. In fact, in recent years there has been a growing interest in amplitude calculations applied to classical gravity. And the benefit is not only on the technical side, but also in obtaining a deeper knowledge that field theory provides for long-standing problems such as interpretation of infrared divergences [79] and the connection between effective potential and scattering angle [80, 81, 82].
The use of effective field theory methods to classical gravity were investigated in a pioneering way by [83], see also [84, 85, 86, 87, 88, 89, 79, 90, 91]. They have been successfully applied to the general problem of two-body relativistic dynamics in the post-Newtonian (PN) approximation up to the fourth perturbative order, with partial results in the fifth perturbative order [92, 93]. The fourth Post-Minkowskian (PM) Order was studied in Ref. [94], and, within amplitude-based methods, in Refs. [41, 42]. In particular, in [95] the connection between the classical two-body gravitational problem and the calculation of quantum field theory amplitudes was highlighted, showing how the evaluation of the effective potential of $n$-th order in PN approximation can be interpreted in terms of a calculation of quantum self-energy in diagrams of massless particles with multiple self-interactions. Solving the equations iteratively in perturbative orders, as usually done in PN approximation, is reminiscent of the application of perturbative methods in quantum field theory, where scattering amplitude calculations are traditionally organized in Feynman diagrams.
In this context, we have witnessed a particularly intense flurry of activities about the gravitational interaction of spinning particles [96, 97, 98, 62, 55, 99, 72, 100, 74, 101, 102, 69, 103, 104, 105, 106, 107, 108, 109, 110, 73, 111, 112, 113, 114, 115, 116, 71, 75, 78, 76, 77, 117, 118, 119, 120, 70, 121, 122, 123, 124, 125, 126, 127, 128, 129]  [^4] For a much larger body of older and current research please see the references within [97, 98, 62, 55, 99, 72, 100, 74, 101, 102, 69, 103, 104, 105, 106, 107, 108, 109, 110, 73, 111, 112, 113, 114, 115, 116, 71, 75, 78, 76, 77, 117, 118, 119, 120, 70, 121, 122, 123, 124, 125, 126, 127, 128, 129].. Indeed, gravitational dynamics of spinning objects has been a subject of vigorous study for several years [130, 131, 132, 133] in both PN and PM approaches. Since scattering amplitudes are envisaged as useful means to incorporate black-hole dynamics, the expectation is that gravitational dynamics of Kerr black holes can be encoded in the dynamics described by a set of higher-dimension operators with presumably known (or to be inferred) Wilson coefficients. Early work suggested that the amplitudes involving massive low-spin states and gravitons could be used to calculate the lowest orders in multipolar spin expansion from black-hole observables, such as scattering angles and gravitational potentials. In fact, conservative observables for Kerr black holes have been studied in a variety of different frameworks, including higher-spin effective field theories and worldline field theories [71, 120, 134, 135, 136].
@@
An outstanding problem emerging in this regard is the evaluation of the gravitational Compton scattering amplitude – the opposite-helicity version is a key ingredient in the evaluation of the conservative potential of compact binary systems with spin [116, 71, 75, 137, 138] and also waveform calculations with spin [129, 128, 139, 140]. Previous calculations in the literature have shown how challenging this task can be [141, 142]  [^5] Opposite-helicity Compton amplitude built from BCFW recursion relations have unphysical poles because this procedure is no longer predictive in the case of higher spins – the product of BCFW shifted three-point amplitudes does not approach zero fast enough as the shift parameter goes to infinity [102]. Nevertheless, we remark that if one starts with a local Lagrangian the amplitudes calculated from it using standard Feynman rules are expected to be manifestly local; see Ref. [116] for a calculation for the case of massive spin 3 electromagnetic Compton.. In any case, subsequent research efforts have successfully navigated initial obstructions and by now important ongoing research is being conducted to address this problem [104, 102, 72, 100, 103, 105, 106, 107, 108, 109, 127, 143, 144].
This scenario also opens up a great opportunity to investigate string amplitudes. Indeed, one of the most notable features of string theory is the presence of an infinite tower of massive spinning modes in its spectrum, required for consistency. Moreover, the string theory framework allows one to consider the scattering of states of indefinite spin in a systematic manner [145]. However, some difficulties have been highlighted in the literature [146, 147, 71]. Nevertheless, as string theory provides a consistent theory of massive interacting higher-spin particles, one can regard it as a natural framework to look into classical objects which involve spin-to-infinity limits.
@@
Indeed, in this paper we wish to report progress in that direction. First, as a promissing result, we extract from a string-theory based calculation the correct tree-level Compton amplitude for a Kerr black hole for quantum spin two  [^6] Paolo di Vecchia and Francesco Alessio have also obtained the same result through a slightly different calculation. We thank them for sharing their unpublished notes with us.. Then, having succeeded in recovering this result, we propose a candidate Compton amplitude which is free from any spurious poles and valid to any integer quantum spin.
In summary, our claim is: string-theory amplitudes could potentially contain amplitudes that describe compact astrophysical bodies, in the form of specific contributions (among many others) in disguise inside the full amplitude. To extract those hidden contributions, all other contributions must be projected out, and we propose that this can be achieved by imposing the following three constraints, to be explained in detail in this paper:
  • soft momentum transfer;
  • compactification of polarizations, so that the mass level matches the spin of the leading Regge states;
Perhaps our procedure is best justified a posteriori, given the results we obtain. With the simple rules above we manage to obtain sensible field-theory Compton amplitudes which are physically sound and whose classical limit agrees with that of Ref. [106] up to fifth order in the ring radius (in the gravitational case). In fact, we argue that such a reasoning can also be used to extract the correct 3-point root-Kerr amplitude from the corresponding superstring amplitude [145, 146], up to terms of order (mass)-1. Throughout the paper, we employ natural units, with $\hbar=c=1$.
@@ 2.1 Open strings
Let us calculate the Compton amplitude involving two photons and two massive spin-$S$ particles. For simplicity in this paper we will only study leading Regge states of the NS sector of the superstring. We consider the Minkowski metric as $\eta_{\mu\nu}=\textrm{diag}(-,+,+,\cdots,+)$ and Mandelstam variables defined as $s_{ij}=-(k_{i}+k_{j})^{2}$, with $s_{12}=s$, $s_{13}=t$ and $s_{14}=u$. In addition, $k_{1}^{2}=k_{3}^{2}=-m^{2}$. Here $s+t+u=2m^{2}=2(S-1)/\alpha^{\prime}$, and the leading Regge state satisfies $m^{2}\alpha^{\prime}=n=S-1$. The vertex operators required here are the following [145]
@@
See also Ref. [148] for the vertex operators of the open bosonic string. We employ a given representation of the massive polarization tensors $\epsilon_{(S)}$ (they are symmetric, traceless and transverse in all indices) of spin-S particles in which they are written as symmetric tensor products of massive spin-$1$ polarization vectors $\epsilon_{1}$, namely $\epsilon_{(S)}=\bigotimes_{S}\epsilon_{1}$. We should also consider $V_{0}^{(0)}$ – in order to cancel the background ghost charge on a genus $g$ Riemann surface, a $g$-loop superstring amplitude needs an overall superghost charge of $22g-$ 2. Finally, we will take $V_{0}^{(0)}$ to represent the photon operator in the zero picture. Our disk amplitude on the leading Regge trajectory reads
@@
where the factor $e^{-\lambda}$ comes from the Euler-number term in the action. The above bold notation refers to massive spinning particles. For convenience we write the amplitude as
@@
where the dot products refer to the contraction of the higher-spin indices for legs 1 and 3. Henceforth we are suppressing the higher-spin indices for clarity. Moreover, we take $z_{1}\to\infty$, $z_{2}=1$ and $z_{4}=0$, which implies that $z_{3}\in[0,1]$.
@@ 2.2 Closed strings
$M_{4}({\bf 1}^{S},2^{h_{2}\tilde{h}_{2}},{\bf 3}^{S},4^{h_{4}\tilde{h}_{4}})$ $=$ $(-1)^{n}\left(\frac{\pi\alpha^{\prime}}{4}\right)^{-1}\sin\left(\frac{\pi\alpha^{\prime}s}{4}\right)$ (8)
@@
$M^{\alpha\gamma,\beta\delta}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $(-1)^{n}\left(\frac{\pi\alpha^{\prime}}{4}\right)^{-1}\sin\left(\frac{\pi\alpha^{\prime}s}{4}\right)\sin\left(\frac{\pi\alpha^{\prime}u}{4}\right)$ (9)
@@ 3 Electromagnetic Compton amplitudes for all spins
3.1 Soft expansion
Our aim is to extract from the above string results terms potentially relevant for the classical description of spinning compact binary systems. Superstring amplitudes are famously defined in ten dimensions, so we need to perform a compactification down to four dimensions; we follow the prescription given by Ref. [146] and define suitable four-dimensional polarizations and momenta. Moreover, we also isolate the poles corresponding to the mass level $n$; these are connected to the s-channel and the u-channel (since we consider electromagnetic amplitudes, the t-channel is absent). To extract our candidate field-theory amplitudes for generic spins, we begin by considering both momenta associated with the external photons as being soft, $q\to\hbar q$. To follow this procedure, we must conveniently rewrite the amplitude by using that
$\frac{\Gamma(1-s\alpha^{\prime})\Gamma(1-u\alpha^{\prime})}{\Gamma(1+\alpha^{\prime}t-2n)}\frac{1}{\alpha^{\prime}t\,\prod_{k=1}^{2n-1}(\alpha^{\prime}t-k)}$ $=$ $\frac{\Gamma(1+2\alpha^{\prime}\,k_{1}\cdot q_{2})}{(1-s\alpha^{\prime})(2-s\alpha^{\prime})\cdots(n-s\alpha^{\prime})}$ (10)
$\times$ $\frac{\Gamma(1+2\alpha^{\prime}\,k_{1}\cdot q_{4})}{(1-u\alpha^{\prime})(2-u\alpha^{\prime})\cdots(n-u\alpha^{\prime})}$
$\times$ $\frac{1}{\Gamma(1+\alpha^{\prime}t)}$
where the recurrence formula
$\Gamma(z)=\frac{\Gamma(z+n+1)}{z(z+1)\cdots(z+n)}$
as well as the result
$\Gamma(z-n)=(-1)^{n-1}\frac{\Gamma(-z)\Gamma(1+z)}{\Gamma(n+1-z)}$
were used. Observe that keeping only the level $n$ poles corresponds to the leading 1/$1/\hbar$ term in the soft expansion. Then at leading order in a soft expansion we find
$A^{\alpha\beta}_{t}\Bigg{|}_{\textrm{field theory}}\equiv{\cal A}^{\alpha\beta}_{S}=\frac{1}{[\Gamma(n)]^{2}}\frac{{\cal A}_{0}^{\alpha\beta}}{(s\alpha^{\prime}-n)(u\alpha^{\prime}-n)}.$ (11)
where the subscript $S$ in ${\cal A}^{\alpha\beta}_{S}$ denotes the spin of the massive particle.
3.2 First mass level
Before continuing the discussion for generic spins, let us study the lowest-spin massive boson. The polarization tensors are given by $\epsilon_{a_{1}a_{S}}({\bf 1})=\epsilon_{a_{1}}({\bf 1})\epsilon_{a_{S}}({\bf 1})$ and $\epsilon_{c_{1}c_{S}}({\bf 3})=\epsilon_{c_{1}}({\bf 3})\epsilon_{c_{S}}({\bf 3})$. By taking a close look at the corresponding expression, one quickly finds terms that describe the Compton amplitude for scattering of massive spin $1$ particles. To separate such terms from the rest we consider a suitable compactification of the polarization vectors $\epsilon_{a_{1}}({\bf 1})$ and $\epsilon_{c_{1}}({\bf 3})$ – we take these to be in the “fifth dimension” (such as $\epsilon^{d=10}({\bf 1})=(0,0,0,0,1,{\bf 0})$). After performing this compactification, contraction with the polarizations, and a few algebraic manipulations, we arrive at the following expression of the Compton amplitude describing the interaction between photons and massive spin-$1$ particles:
@@ 3.2 First mass level
where $h_{2},h_{4}$ are the helicities of the photons and $f_{i}^{\mu\nu}=2q_{i}^{[\mu}\epsilon^{\nu]}(i)$. To obtain this expression we wrote $\epsilon_{a_{S}}({\bf 1})\epsilon_{c_{S}}({\bf 3})={\cal J}_{a_{S}c_{S}}/2+\epsilon_{(a_{S}}({\bf 1})\epsilon_{c_{S})}({\bf 3})$, where we have defined ${\cal J}^{\alpha\gamma}\equiv-i\epsilon^{\mu}({\bf 1})(M^{\alpha\gamma})_{\mu\nu}\epsilon^{\nu}({\bf 3})$, $(M^{ab})_{c}^{\ d}=2i\delta^{[a}_{\ c}\eta^{b]d}$ being the Lorentz generator in the spin-$1$ representation. Furthermore:
@@
Up to a numerical overall factor, the result (12) perfectly matches the one previously derived by others [100] (except for possible sign differences due to different conventions on momenta).
The spin multipole decomposition of the Compton amplitude given by (12) can be easily rewritten in the massive spinor-helicity basis [151]. For the opposite helicity amplitude, $h_{2}=+,h_{4}=-$, one finds that
$A_{S}({\bf 1},2^{+},{\bf 3},4^{-})$ $=$ $-\frac{2ie^{2}}{m^{2S}}\frac{\langle 4|{\bf 1}|2\bigr{]}^{2}}{(s-m^{2})(u-m^{2})}\langle{\bf 3}|^{2S}\exp\left(i\frac{q_{4}^{\mu}\bar{\epsilon}_{4}^{\nu}{\cal J}_{1\mu\nu}}{\bar{p}\cdot\bar{\epsilon}_{4}}\right)|{\bf 1}\rangle^{2S}$ (14)
$=$ $-\frac{2ie^{2}}{(s-m^{2})(u-m^{2})}\langle 4|{\bf 1}|2\bigl{]}^{2-2S}\Bigl{(}\langle{\bf 3}4\rangle\bigl{[}{\bf 1}2\bigr{]}+\langle{\bf 1}4\rangle\bigl{[}{\bf 3}2\bigr{]}\Bigr{)}^{2S}\,\,\,\,\,\,\,\,(S\leq 1)$
where $\bar{p}=(p_{1}-p_{3})/2$ and $\bar{\epsilon}_{4}$ is the polarization vector for $q_{4}$ with reference momentum fixed to be $q_{2}$, see also [113]. A similar calculation can be performed for the same-helicity amplitude. Therefore the spin-1 amplitude, computed from a string-theory based calculation, agrees with other known expressions in the literature [141, 142].
Observe from Eq. (10) that $\alpha^{\prime}$ is effectively parameterizing the expansion in the soft photon momenta, since it always appears accompanied by a Mandelstam variable. Nevertheless, to see that the approach of taking directly $\alpha^{\prime}\to 0$ is not, in general, the proper way to obtain the field theory limit which is useful in the present context, let us quote the amplitude for the spin-2 leading Regge state with this prescription. The calculation is rather involved but straightforward. After the dust settles, we find that
@@
where ${\cal J}^{ab}\equiv-\frac{i}{2}\epsilon^{c_{1}c_{2}}({\bf 1}){(M^{ab})_{c_{1}c_{2}}}^{d_{1}d_{2}}\epsilon_{d_{1}d_{2}}({\bf 3})$, ${(M^{ab})_{c(2)}}^{d(2)}=4i\delta_{(c_{1}}^{[a}\eta^{b](d_{1}}\delta_{c_{2})}^{d_{2})}$ being the Lorentz generator in the spin-$2$ representation. This amplitude in the classical limit will produce contributions up to the quadrupole term. However, for a spin-$S$ quantum particle, the classical counterpart should produce an expansion up to the $2S$-th power of the classical spin. Besides, strictly speaking, the $\alpha^{\prime}\to 0$ amounts to projecting out all massive spinning particles from the spectrum. These are clearly indications that the straightforward limit $\alpha^{\prime}\to 0$ is not adequate to our purposes.
3.3 All mass levels
Given the precedent successful computation of the amplitude for the massive spin-$1$ case, Eq. (12), we now undertake the task of investigating the general case, with the polarization tensors given by $\epsilon_{a_{1}\cdots a_{S}}({\bf 1})=\epsilon_{a_{1}}({\bf 1})\ldots\epsilon_{a_{n}}({\bf 1})\epsilon_{a_{S}}({\bf 1})$ and $\epsilon_{c_{1}\cdots c_{S}}({\bf 3})=\epsilon_{c_{1}}({\bf 3})\ldots\epsilon_{c_{n}}({\bf 3})\epsilon_{c_{S}}({\bf 3})$. Following previous prescriptions, we consider the compactification of two massive polarization vectors, namely $\epsilon_{a_{n}}({\bf 1})$ and $\epsilon_{c_{n}}({\bf 3})$. Furthermore, we consider only terms with $k=l$ in the double summations in the full expression for ${\cal A}_{0}^{\alpha\beta}$ (see appendix) as these are the ones that have explicit inverse propagators in the numerators. Our candidate electromagnetic Compton amplitude takes the form